# Slant sums of quiver gauge theories

### v1

@@ 5.3 Torus action and fixed points on Coulomb branches
The space $\mathcal{R}$ maps naturally to the affine Grassmannian $\operatorname{Gr}_{G}=G_{\mathcal{K}}/G_{\mathcal{O}}$. Connected components of $\operatorname{Gr}_{G}$ are labeled by the lattice $\pi_{1}(G)$. So, $\mathbb{C}[\mathcal{M}_{C}]$ admits a grading by $\pi_{1}(G)$, and hence, the variety $\mathcal{M}_{C}$ admits a (Hamiltonian) action of the torus $\pi_{1}(G)^{\wedge}$, where by $\bullet^{\wedge}$ we mean the Pontryagin dual (see [8, Section 3(v)] for the details). The cohomological grading on $\mathbb{C}[\mathcal{M}_{C}]=H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ defines the action of $\mathbb{C}^{\times}$ on $\mathcal{M}_{C}$.
A choice of the stability parameter $\theta\colon G\rightarrow\mathbb{C}^{\times}$ defines a cocharacter $\mathbb{C}^{\times}\rightarrow\pi_{1}(G)^{\wedge}$. Abusing notations, we will denote this cocharacter by $\theta$. In [6, Conjecture 3.25(1)] the authors conjecture that the set $\mathcal{M}_{C}^{\theta(\mathbb{C}^{\times})}(\mathbb{C})$ of $\mathbb{C}$-points of the schematic fixed points of $\mathcal{M}_{C}$ is either empty or consists of one point. From now on, we denote $\mathcal{M}_{C}^{\theta(\mathbb{C}^{\times})}$ simply by $\mathcal{M}_{C}^{\theta}$.
@@ 5.4 Mirror symmetry and some conjectures
In this section we formulate various conjectures relating Higgs and Coulomb branches for quiver theories.
3d mirror symmetry is a nontrivial relation between Higgs and Coulomb branches for $(G,{\bf{N}})$. From now on, we assume that our theory is a quiver gauge theory for a quiver $Q$ without loops, and we denote the smooth symplectic variety $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ by $\widetilde{\mathcal{M}}_{H}$ (recall that in many cases $\widetilde{\mathcal{M}}_{H}$ is a symplectic resolution of the Higgs branch $\mathcal{M}_{H}$).
One mirror symmetry statement is the Hikita conjecture [20] predicting an identification of (graded) algebras:
@@
So, it is natural to expect that $\mathcal{M}_{C}^{\theta}$ is nonempty if and only if $\widetilde{\mathcal{M}}_{H}$ is nonempty.
We start with the following standard lemma.
Lemma 5.1.
Let $\pi\colon X\rightarrow Y$ be a conical symplectic resolution. Assume that $H^{*}(X)=\mathbb{C}$. Then $X$ is isomorphic to $\mathbb{A}^{2k}$, where $2k=\operatorname{dim}X$.
Proof.
Let $p\in Y$ be the unique $\mathbb{C}^{\times}$-fixed point of $Y$ (recall that $Y$ contracts to $p$ via the action of $\mathbb{C}^{\times}$). The $\mathbb{C}^{\times}$-action contracts $X$ to $X^{\mathbb{C}^{\times}}=\pi^{-1}(p)^{\mathbb{C}^{\times}}$, so $H^{*}(X)=H^{*}(\pi^{-1}(p)^{\mathbb{C}^{\times}})$. Note now that $\pi^{-1}(p)^{\mathbb{C}^{\times}}$ is smooth and proper, so $H^{\mathrm{top}}(\pi^{-1}(p)^{\mathbb{C}^{\times}})=\mathbb{C}$. It follows that $\pi^{-1}(p)$ consists of one point. We conclude that $X$ contracts to one point under the action of $\mathbb{C}^{\times}$, hence, by the Bialynicki-Birula theorem (see [4]), $X$ must be isomorphic to the affine space. ∎
We make the following conjecture:
@@
The variety $\widetilde{\mathcal{M}}_{H}$ is a point if and only if $\mathcal{M}_{C}^{\theta}$ consists of a single nonsingular point.
Let us explain our motivation for this conjecture. If $p^{!}\in\mathcal{M}_{C}$ is nonsingular, then $\mathbb{C}[\mathcal{M}_{C}^{\theta}]$ must be equal to $\mathbb{C}$. Assuming (12), we deduce that $H^{*}(\widetilde{\mathcal{M}}_{H})=\mathbb{C}$. Then Lemma 5.1 implies that $\widetilde{\mathcal{M}}_{H}$ is isomorphic to $\mathbb{A}^{2k}$. We expect that for quivers $Q$ without edge loops, the dimension of $\widetilde{\mathcal{M}}_{H}$ is actually equal to zero whenever $p^{!}\in\mathcal{M}_{C}$ is nonsingular.
Recall that Conjecture 1.5 gives a formula for the tangent space $T_{p^{!}}\mathcal{M}_{C}$ when $\widetilde{\mathcal{M}}_{H}$ is a point. Actually we expect more to be true: this formula should allow one to compute the character of the tangent space for an arbitrary nonsingular fixed point of a partially resolved Coulomb branch. Let us explain the details.
Let $\mathsf{A}$ be a torus acting on $\bf{N}$} commuting with $G$ and let $\nu\colon\mathbb{C}^{\times}\rightarrow\mathsf{A}$ be a cocharacter. As in [8, Section 3(ix)], it defines a partial resolution
@@
where the disjoint union is taken over all conjugacy classes of cocharacters $\gamma\colon\mathbb{C}^{\times}\rightarrow G$. Moreover, every $(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ corresponds to some quiver, hence, each $\widetilde{\mathcal{M}}_{H}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ itself is a Nakajima quiver variety.
@@
The following conjecture is due to Justin Hilburn (see [23, Conjecture 1.6]).
@@
$j\colon\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})\hookrightarrow\widetilde{\mathcal{M}}_{C,\nu}.$
@@
Assuming both Conjectures 5.2 and 5.3 hold, one obtains the following corollary.
@@
Recall now that Conjecture 1.5 describes the tangent space $T_{p^{!}}\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$. So, combining it with Conjecture 5.3, we should obtain a conjectural description of the tangent space to an nonsingular $\theta$-fixed point of $\widetilde{\mathcal{M}}_{C,\nu}$.
@@
We expect that there should be a bijection between nonsingular $\theta$-fixed points of $\widetilde{\mathcal{M}}_{C,\nu}$ and isolated $\nu$-fixed points on $\widetilde{\mathcal{M}}_{H}$. If that is the case, then a tangent space to arbitrary non-singular $\theta$-fixed point should have a description as above. Note that, more generally, it is reasonable to expect that there should be a bijection between the set $\widetilde{\mathcal{M}}_{C,\nu}^{\theta}(\mathbb{C})$ and the set $\operatorname{Comp}(\widetilde{\mathcal{M}}_{H}^{\nu})$ of components of the $\nu$-fixed points of $\widetilde{\mathcal{M}}_{H}^{\nu}$ (this is another incarnation of the equivariant Hikita conjecture).
@@
Note that we are not assuming that the torus $\mathsf{A}$ comes from the action of $\prod_{i}GL(W_{i})$ on $\bf{N}$}. If this is the case, then $\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ is the product of Coulomb branches for the same quiver $Q$.
Let us finish this section by mentioning that whenever Conjecture 1.5 is proven, we obtain an explicit formula for a normalized character of the (unique) irreducible module in the category $\mathcal{O}$ for $\mathcal{A}$. Namely, assume that Conjecture 1.5 holds for some $Q,\mathsf{v},\mathsf{w}$. Fix $x\in\mathfrak{a}$ and let
$\mathcal{A}_{x}:=H_{*}^{(G_{\mathcal{O}}\rtimes\mathbb{C}^{\times})\times\mathsf{A}}(\mathcal{R})_{(x,1)}$
be the corresponding quantized Coulomb branch algebra. The cocharacter $\theta$ determines the category $\mathcal{O}_{\theta}(\mathcal{A}_{x})$ as in [27, Section 3.4]. By, [27, Lemma 3.27], irreducible objects of this category are in bijection with irreducible modules over the (finite-dimensional) algebra $\mathsf{C}_{\theta}(\mathcal{A}_{x})$ called Cartan subquotient or $B$-algebra and defined as follows:
$\mathsf{C}_{\theta}(\mathcal{A}_{x}):=\mathcal{A}_{x,0}/\sum_{i<0}\mathcal{A}_{x,i}\mathcal{A}_{x,-i},$
where index $i$ corresponds to the $\mathbb{Z}$-grading on $\mathcal{A}_{x}$ induced by the cocharacter $\theta$. It is known (see [27, Lemma 3.24]) that there exists a filtration on $\mathsf{C}_{\theta}(\mathcal{A}_{x})$ such that the associated graded $\operatorname{gr}\mathsf{C}_{\theta}(\mathcal{A}_{x})$ is a quotient of $\mathbb{C}[\mathcal{M}_{C}^{\theta}]$. Note now that $\mathbb{C}[\mathcal{M}_{C}^{\theta}]\simeq\mathbb{C}$, since $p^{!}=\mathcal{M}_{C}^{\theta}(\mathbb{C})$ is a nonsingular point of $\mathcal{M}_{C}$. So, we conclude that the category $\mathcal{O}_{\theta}(\mathcal{A}_{x})$ has a unique irreducible object $L=L(p^{!})$. Let $\widetilde{\operatorname{ch}}(L)$ be the $\mathsf{A}_{Q}$-character of $L$ normalized in such a way that it starts with $1$.
Conjecture 5.7.
For $\theta>0$ we have:
$\widetilde{\operatorname{ch}}(L)=\prod_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\frac{1}{(1-e^{\alpha})^{\langle\alpha,\mu\rangle}}.$ (14)

### v2

@@ 5.3 Torus action and fixed points on Coulomb branches
The space $\mathcal{R}$ maps naturally to the affine Grassmannian $\operatorname{Gr}_{G}=G_{\mathcal{K}}/G_{\mathcal{O}}$. Connected components of $\operatorname{Gr}_{G}$ are labeled by the lattice $\pi_{1}(G)$. So, $\mathbb{C}[\mathcal{M}_{C}]$ admits a grading by $\pi_{1}(G)$, and hence, the variety $\mathcal{M}_{C}$ admits a (Hamiltonian) action of the torus $\pi_{1}(G)^{\wedge}$, where by $\bullet^{\wedge}$ we mean the Pontryagin dual (see [undefj, Section 3(v)] for the details). The cohomological grading on $\mathbb{C}[\mathcal{M}_{C}]=H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ defines the action of $\mathbb{C}^{\times}$ on $\mathcal{M}_{C}$.
A choice of the stability parameter $\theta\colon G\rightarrow\mathbb{C}^{\times}$ defines a cocharacter $\mathbb{C}^{\times}\rightarrow\pi_{1}(G)^{\wedge}$. Abusing notations, we will denote this cocharacter by $\theta$. In [undefh, Conjecture 3.25(1)] the authors conjecture that the set $\mathcal{M}_{C}^{\theta(\mathbb{C}^{\times})}(\mathbb{C})$ of $\mathbb{C}$-points of the schematic fixed points of $\mathcal{M}_{C}$ is either empty or consists of one point. From now on, we denote $\mathcal{M}_{C}^{\theta(\mathbb{C}^{\times})}$ simply by $\mathcal{M}_{C}^{\theta}$.
@@ 5.4 Mirror symmetry and some conjectures
In this section we formulate various conjectures relating Higgs and Coulomb branches for quiver gauge theories.
5.4.1 Coulomb branches with non-singular torus fixed points
3d mirror symmetry is a nontrivial relation between Higgs and Coulomb branches for $(G,{\bf{N}})$. From now on, we assume that our theory is a quiver gauge theory for a quiver $Q$ without edge loops, and we denote by $\mathfrak{g}=\mathfrak{g}_{Q}$ the corresponding Kac-Moody Lie algebra. We will denote the smooth symplectic variety $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ by $\widetilde{\mathcal{M}}_{H}$ (recall that in many cases $\widetilde{\mathcal{M}}_{H}$ is a symplectic resolution of the Higgs branch $\mathcal{M}_{H}$).
For $i\in Q_{0}$ we will denote by $\omega_{i}^{\vee},\alpha_{i}^{\vee}$ the corresponding fundamental coweight and simple coroot for $\mathfrak{g}$. Given $\mathsf{v},\mathsf{w}\in\mathbb{N}^{Q_{0}}$, set:
$\lambda=\sum_{i}\mathsf{w}_{i}\omega_{i}^{\vee},~\mu=\lambda-\sum_{i}\mathsf{v}_{i}\alpha_{i}^{\vee}.$ (13)
One mirror symmetry statement is the Hikita conjecture [undefy] predicting an identification of (graded) algebras:
@@ 5.4.1 Coulomb branches with non-singular torus fixed points
If $\widetilde{\mathcal{M}}_{H}$ is a point then, assuming (14), we deduce that $\mathbb{C}[\mathcal{M}_{C}^{\theta}]=\mathbb{C}$ that suggests that $p^{!}\in\mathcal{M}_{C}$ should be nonsingular. It follows from Remark 7.5 below that $\widetilde{\mathcal{M}}_{H}$ is a point iff $\mu\in W\lambda$. So, we arrive to the following Conjecture that can be formulated without referring to $\widetilde{\mathcal{M}}_{H}$ and so has a chance to be valid for arbitrary symmetrizable quivers $Q$ (where $\mathcal{M}_{C}$ is defined as in [undefap]).
@@
We have $\mathcal{M}_{C}^{\theta}=\{p^{!}\}$ and $p^{!}\in\mathcal{M}_{C}$ is nonsingular if $\mu\in W\lambda$.
Remark 5.2.
It’s natural to expect that the implication of Conjecture 5.1 also holds in the opposite direction. Namely if $p^{!}\in\mathcal{M}_{C}$ is nonsingular then $\mu\in W\lambda$. Note that, assuming (14) one would see that $H^{*}(\widetilde{\mathcal{M}}_{H})=\mathbb{C}$. It’s not hard to show that $H^{*}(\widetilde{\mathcal{M}}_{H})=\mathbb{C}$ forces $\widetilde{\mathcal{M}}_{H}$ to be isomorphic to the affine space.
5.4.2 Tangent spaces to nonsingular fixed points on partial resolutions
Recall that Conjecture 1.7 gives a formula for the tangent space $T_{p^{!}}\mathcal{M}_{C}$ when $\mu\in W\lambda$. Actually we expect more to be true: this formula should allow one to compute the character of the tangent space for an arbitrary nonsingular fixed point of a partially resolved Coulomb branch. Let us explain the details.
Let $\mathsf{A}$ be a torus acting on $\bf{N}$} commuting with $G$ and let $\nu\colon\mathbb{C}^{\times}\rightarrow\mathsf{A}$ be a cocharacter. As in [undefj, Section 3(ix)], it defines a partial resolution
@@ 5.4.2 Tangent spaces to nonsingular fixed points on partial resolutions
where the disjoint union is taken over all conjugacy classes of cocharacters $\gamma\colon\mathbb{C}^{\times}\rightarrow G$. Moreover, every $(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ corresponds to some quiver, hence, each $\widetilde{\mathcal{M}}_{H}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ itself is a Nakajima quiver variety, see, for example [undefv, Proposition 2.13].
@@
The following conjecture is due to Justin Hilburn (see [undefae, Conjecture 1.6]).
@@
$j\colon\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})\hookrightarrow\widetilde{\mathcal{M}}_{C,\nu}.$ (16)
@@
Assuming both Conjectures 5.1 and 5.3 hold, one obtains the following corollary.
@@
Recall now that Conjecture 1.7 describes the tangent space $T_{p^{!}}\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$. So, combining it with Conjecture 5.3, we should obtain a conjectural description of the tangent space to a nonsingular $\theta$-fixed point of $\widetilde{\mathcal{M}}_{C,\nu}$. We expect that there should be a bijection between nonsingular $\theta$-fixed points of $\widetilde{\mathcal{M}}_{C,\nu}$ and isolated $\nu$-fixed points on $\widetilde{\mathcal{M}}_{H}$. If that is the case, then a tangent space to arbitrary non-singular $\theta$-fixed point of $\widetilde{\mathcal{M}}_{C,\nu}$ should have a description as above.
@@
Note that we are not assuming that the torus $\mathsf{A}$ arises from the action of $\prod_{i}GL(W_{i})$ on $\bf{N}$} (i.e., that it is a framing torus). If it does, then $\mathcal{M}_{C}(Z_{G}(\gamma),{\bf{N}}^{(\gamma,\nu)})$ is a product of Coulomb branches for the same quiver $Q$. Note also, that for our purposes it suffices to have a much weaker version of Conjecture 16, namely that the completions of (16) at the corresponding fixed points are $\pi_{1}(G)^{\wedge}$-isomorphic. One possible way to prove this is by showing that this holds after deforming $\widetilde{\mathcal{M}}_{C,\nu}$.
6 Irreducible modules over some Coulomb branches
Assume that $Q,\mathsf{v},\mathsf{w}$ are such that $\mathcal{M}_{C}^{\theta}=\{p\}$ and $p\in\mathcal{M}_{C}$ is nonsingular. Recall that conjecturally this holds iff $\mu\in W\lambda$, where $\lambda,\mu$ are constructed from ${\mathsf{v}},\mathsf{w}$ as in (2).
Fix $x\in\mathfrak{a}_{Q}$ and let
$\mathcal{A}_{x}:=H_{*}^{(G_{\mathcal{O}}\rtimes\mathbb{C}^{\times})\times\mathsf{A}_{Q}}(\mathcal{R})_{(x,1)}$
be the corresponding quantized Coulomb branch algebra. To simplify the notation, we denote $\mathcal{A}_{x}$ simply by $\mathcal{A}$. Cocharacter $\theta$ determines the category $\mathcal{O}_{\theta}(\mathcal{A})$ as in [undefak, Section 3.4].
It’s not hard to see that the category $\mathcal{O}_{\theta}(\mathcal{A})$ contains one simple object (to be denoted by $L$, see Section 6.1 below for the details). The goal of this section is to show that the normalized $\mathsf{a}_{Q}$-character of $L$ coincides with the character of $T_{p}^{-}\mathcal{M}_{C}$, in particular, it (and it’s $\hbar^{!}$-deformed version) can be read off from the equation (5) whenever Conjecture 1.7 holds.
6.1 Category $\mathcal{O}$
Recall that the cocharacter $\theta$ defines the category $\mathcal{O}_{\theta}(\mathcal{A})$. By, [undefak, Lemma 3.27], irreducible objects of this category are in bijection with irreducible modules over the (finite-dimensional) algebra $\mathsf{C}_{\theta}(\mathcal{A})$ called Cartan subquotient or $B$-algebra and defined as follows:
$\mathsf{C}_{\theta}(\mathcal{A}):=\mathcal{A}_{0}/\sum_{i<0}\mathcal{A}_{i}\mathcal{A}_{-i},$
where index $i$ corresponds to the $\mathbb{Z}$-grading on $\mathcal{A}$ induced by the cocharacter $\theta$. It follows from Lemma 6.7 below that $\mathsf{C}_{\theta}(\mathcal{A})=\mathbb{C}$, so $\mathcal{O}_{\theta}(\mathcal{A})$ has the unique irreducible object to be denoted $L$.
@@ 6.1 Category $\mathcal{O}$
Alternatively, assuming $\mu\in W\lambda$ one can use [undefag, Theorem 9.14] to see that $\mathcal{O}_{\theta}(\mathcal{A})$ has a unique irreducible object. We are to grateful to Joel Kamnitzer for pointing this out.
We have an embedding $\mathsf{a}_{Q}\subset\mathcal{A}$ inducing the action $\mathsf{a}_{Q}\curvearrowright L$. By normalized $\mathsf{a}_{Q}$-character of $L$ we mean the $\mathsf{a}_{Q}$-character of $L$ normalized in such a way that it starts with $1$.
6.2 Completions and a formal version of $L$
The algebra $\mathcal{A}$ is $\mathbb{Z}$-filtered and the filtration is complete and separated. For $k\in\mathbb{Z}$ let us denote the $k$’th filtration term by $F^{k}\mathcal{A}$. Algebra $\mathcal{A}$ is also $\mathbb{Z}$-graded, where the grading is induced by the $\theta(\mathbb{C}^{\times})$-action. For $i\in\mathbb{Z}$ we denote by $\mathcal{A}_{i}$ the corresponding graded component. Set:
$\widehat{\mathcal{A}}=\underset{\underset{k\rightarrow-\infty}{\longleftarrow}}{\operatorname{lim}}\,\mathcal{A}/F^{k}\mathcal{A},$
algebra $\widehat{\mathcal{A}}$ is the completion of $\mathcal{A}$ w.r.t. the filtration $F^{\bullet}\mathcal{A}$. Filtration $F^{\bullet}$ on $\mathcal{A}$ induces the filtration on $\widehat{\mathcal{A}}$ to be denoted by the same symbol. We have $\operatorname{gr}\widehat{\mathcal{A}}=\operatorname{gr}\mathcal{A}=\mathbb{C}[\mathcal{M}_{C}]$.
For a $\mathbb{Z}$-filtered vector space $F^{\bullet}V$, we set $\operatorname{R}_{\hbar}(V):=\bigoplus_{k\in\mathbb{Z}}(F^{k}V)\hbar^{k}\subset V[\hbar^{\pm 1}]$ and set $\widehat{\operatorname{R}}_{\hbar}(V):=\underset{\underset{k\rightarrow+\infty}{\longleftarrow}}{\operatorname{lim}}\,\operatorname{R}_{\hbar}(V)/(\hbar^{k})$.
Set:
[…]
We have $\widehat{\mathcal{A}}_{\hbar}/(\hbar)=\operatorname{gr}\widehat{\mathcal{A}}=\mathbb{C}[\mathcal{M}_{C}]$. Let $\mathfrak{m}_{p}\subset\mathbb{C}[\mathcal{M}_{C}]$ be the maximal ideal of $p\in\mathcal{M}_{C}$ and let $\mathfrak{m}_{p,\hbar}\subset\widehat{\mathcal{A}}_{\hbar}$ be its preimage. Set
$\widehat{\mathcal{A}}_{\hbar,\widehat{p}}:=\underset{\underset{n\rightarrow+\infty}{\longleftarrow}}{\operatorname{lim}}\,\widehat{\mathcal{A}}_{\hbar}/\mathfrak{m}_{p,\hbar}^{n}.$
Let’s describe the algebra $\widehat{\mathcal{A}}_{\hbar,\widehat{p}}$. First of all recall that $\mathcal{A}$ admits an action of $\mathsf{A}_{Q}$, it induces the action on $\widehat{\mathcal{A}}_{\hbar,\widehat{p}}$.
For a vector space $V$ equipped with a skew-symmetric pairing (,)(\,,\,) let $\widehat{\mathbb{W}}_{\hbar}(V)$ be the formal Weyl algebra over $\mathbb{C}[[\hbar]]$ defined as follows:
$\widehat{\mathbb{W}}_{\hbar}(V)=\mathbb{C}[[V,\hbar]]$
subject to relations
$\hbar,V]=0,[a,b]=\hbar(a,b),~a,b\in V.$ (17)
We will be interested in the vector space $V=T^{*}_{p}\mathcal{M}_{C}$ with a pairing (,)(\,,\,) induced by the sympletic form on $\mathcal{M}_{C}^{\mathrm{reg}}$. Note that the torus $\mathsf{A}_{Q}$ acts on $T^{*}_{p}\mathcal{M}_{C}$ inducing the action on $\widehat{\mathbb{W}}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})$ (by algebra automorphisms).
Lemma 6.2.
There exists a $\mathsf{A}_{Q}$-equivariant isomorphism of algebras over $\mathbb{C}[[\hbar]]$:
$\widehat{\mathbb{W}}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})\simeq\widehat{\mathcal{A}}_{\hbar,\widehat{p}}$.
Proof.
Algebra $\widehat{\mathcal{A}}_{\hbar,\widehat{p}}$ is a formal quantization of $\mathbb{C}[\widehat{p}]=\underset{\underset{n\rightarrow+\infty}{\longleftarrow}}{\operatorname{lim}}\,\mathbb{C}[\mathcal{M}_{C}]/\mathfrak{m}_{p}^{n}$. Note that $\mathbb{C}[\widehat{p}$] as a $\mathsf{A}_{Q}$-graded Poisson algebra is isomorphic to $\mathbb{C}[[T_{p}\mathcal{M}_{C}]]$. Any formal quantization of $\mathbb{C}[[T_{p}\mathcal{M}_{C}]]$ is isomorphic to $\widehat{\mathbb{W}}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})$ and this isomorphism can be made ${\mathsf{A}}_{Q}$-equivariant (see [undefc, Lemma 1.5] and [undefg, Lemma 5.2]). ∎
Recall now that whenever a torus $\mathbb{T}$ acts on some vector space $R$, we can consider a vector subspace $R^{\mathrm{fin}}\subset R$ generated by the eigenvectors of $\mathbb{T}\curvearrowright R$. We will call $R^{\mathrm{fin}}\subset R$ the subspace of $\mathbb{T}$-finite vectors.
We define:
$\mathbb{W}_{\hbar}(T^{*}_{p}\mathcal{M}_{C}):=\widehat{\mathbb{W}}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})^{\mathrm{fin}},$
where $\bullet^{\mathrm{fin}}$ is taken with respect to the torus $\mathsf{A}_{Q}$.